# FELA: A Multi-Agent Evolutionary System for Feature Engineering of Industrial Event Log Data

### v2

@@ V-B1 Datasets and Feature Engineering Task
In the experiment, three real-world datasets are adopted for evaluation, including Diabetes Health Indicator Dataset (Dia) [55], Tabao Conversion Prediction Data (Taobao) [56], and the User Churn Data in Tencent Game Platform (Tencent). The basic information of the three datasets is summarized in Table I, where Dia is standard tabular data while the others are event log data with real industrial complexity.
@@
  • Taobao Dataset: This industrial log dataset contains 176,773 users, 9599 items, and 420,721 ad records over 7 days, where each record is labeled with a binary ad impression. The data is organized across five relational tables containing basic identifiers, item profiles, user demographics, contextual signals, and shop reputation metrics, which constitute 31 features for each record. The feature engineering functions jointly process the entire log data and output the features for each user-ad pair. Following standard temporal evaluation protocols, the dataset is chronologically split with the initial 6 days for training and the final day for testing, effectively simulating real-world conversion prediction scenarios where models forecast future conversion behavior based on learned historical patterns.
  • Tencent Dataset: This dataset is collected from a Massively Multiplayer Online (MMO) game in Tencent for a few days, where the players are sampled from the same user profile to satisfy the i.i.d assumption. Each player is associated with a churn label, indicating whether the user will churn in the next day. In our scenario, each player can have multiple game plays in a day, which results in multiple event log instances. Furthermore, each player in the same game can have multiple teammates, and if the player has more than one teammate, we will separate the records into different rows. Therefore, the instances in such event log data do not follow the i.i.d assumption as mentioned in the introduction session. The feature engineering functions jointly process the entire log data and output the features for each player. The downstream churn prediction is thus working on player level. For evaluation, a temporal split is adopted: the former day of data are used for training, and the logs from the latter day are held out for testing.
@@ V-B2 Baselines
We evaluate the performance of our proposed FELA system against three baselines: (1) raw features, (2) FeatLLM [57], and (3) LLM-FE [12], where FeatLLM and LLM-FE are state-of-the-art free coding methods. To ensure a fair comparison, all LLM-based methods, including the baselines and our proposed FELA, utilize the DeepSeek-V3 model via the DeepSeek API [58].
@@ V-B3 Metrics
The feature engineering for FELA and the LLM-FE baseline is optimized using the AUC-score, which we denote as the target metric $\mathcal{R}$ in Equation (1). The generated features of all methods are then evaluated using a Random Forest classifier by default, with performance reported across five standard metrics: Accuracy, Precision, Recall, F1-score, and AUC.
TABLE II: Performance comparison on Dia dataset.
Method Acc Prec Recall F1 AUC
Raw Features 0.859 0.489 0.166 0.247 0.797
@@
V-C Feature Engineering Performance (RQ1)
Table II presents the feature engineering performance of FELA and baseline methods on the tabular Dia dataset. The proposed FELA achieves superior performance in accuracy, precision, and AUC, demonstrating its effectiveness in handling structured data. Notably, FELA improves precision from 0.514 to 0.552 and AUC from 0.802 to 0.812 compared to the runner-up method LLM-FE, indicating its capability to generate high-quality discriminative features for tabular data analysis.
On the Taobao conversion prediction dataset, FELA maintains its competitive advantage as shown in Table III. Our method obtains the highest scores in accuracy, precision, F1-score, and AUC, with a notable AUC improvement from 0.641 to 0.653 over LLM-FE. This consistent superiority across multiple metrics confirms FELA’s robustness in processing complex industrial log data with relational characteristics.
The evaluation on the Tencent user churn dataset further validates FELA’s superiority on complex industrial log data. As summarized in Table IV, FELA outperforms all baselines in accuracy, precision, F1-score, and AUC, achieving a significant AUC boost from 0.686 to 0.701 compared to LLM-FE. These results across three distinct domains demonstrate that FELA consistently generates semantically meaningful features that enhance predictive performance in diverse real-world industrial scenarios.
TABLE III: Performance comparison on Taobao dataset.
Method Acc Prec Recall F1 AUC
Raw Features 0.630 0.0269 0.580 0.0514 0.630
FeatLLM 0.645 0.026 0.536 0.0496 0.631
@@ V-C Feature Engineering Performance (RQ1)
TABLE IV: Performance comparison on Tencent dataset.
Method Acc Prec Recall F1 AUC
Raw Features 0.668 0.576 0.385 0.461 0.683
@@
Figure 7: AUC comparison across different classifiers.
@@ V-D Steerable Feature Generation (RQ2)
We begin by analyzing the complexity of the generated feature engineering code on the Taobao dataset, as illustrated in Figure 8. After 16 hours of execution, the proposed FELA framework generates 22 features derived from 5 distinct feature ideas, achieving an AUC score of 0.653, an improvement of 0.023 over the raw features. This demonstrates FELA’s capability to efficiently produce a substantial number of high-quality features. In contrast, the FeatLLM baseline generates 20 new features within 2 minutes, yet yields negligible AUC improvement compared to raw features. Meanwhile, LLM-FE produces 110 code snippets within 30 minutes, with the best-performing code improving AUC by 0.011. However, the optimal LLM-FE code only creates only 1 new feature and lacks a continuous evolution mechanism, ultimately limiting its performance potential.
@@
Figure 9: Correlations of selected ideas & features in FELA.
We further investigate the feature correlation patterns in FELA by analyzing the Pearson correlation heatmap of the aforementioned 22 features, as visualized in Figure 9. The analysis reveals that features derived from the same idea exhibit stronger intra-idea correlation. This phenomenon aligns with expectations, as features originating from the same semantic insight naturally share underlying conceptual foundations. Conversely, inter-idea feature correlations remain relatively low, indicating that FELA successfully explores diverse regions of the feature space. This structured exploration strategy demonstrates FELA’s capability to conduct steerable feature generation while maintaining semantic interpretability. The framework effectively balances between exploiting productive semantic directions and exploring novel feature combinations, resulting in both discriminative power and explanatory value. In Figure 10, we visualize the sampled knowledge database after evolution on the Taobao dataset. We can easily trace the evolution of knowledge base structure, as well as understand the evolutionary trajectories between the ideas and associated features.
Figure 10: A snapshot of the knowledge database during evolution. It is a super set of the selected features. Green denotes synthesized ideas. Purple denotes new ideas. Red denotes original ideas. Light blue denotes derived features.
@@ V-E Compatibility to Various Machine Learning Models (RQ3)
We investigate the applicability of the proposed FELA over various machine learning models, including Random Forest, XGBoost, MLP, and Logistic classifiers. Then, the AUC performance of the proposed FELA and the baselines over the Tencent dataset is shown in Figure 7. It can be observed that the proposed FELA can achieve the best AUC performance over most of the classifiers, where a 0.012 AUC improvement can be achieved on average. Thus, the proposed FELA system can serve as a robust feature engineering solution for various machine learning models.
@@ V-F Abalation Study (RQ4)
We conduct the ablation study in FELA systems to investigate the contribution of its components, including the critic agent, long-short-term memory, and the UCB-based learning algorithm. Then, the AUC in different iterations for the proposed FELA and its ablations is plotted in Figure 11. The ablations without the critic agent and the long-short term memory lead to an obvious AUC decline of around 0.01. Thus, the proposed FELA can benefit from long-short-term memory to exploit the evolution experience. Additionally, the critic agent can enhance the correctness and consistency of the generated idea and code, which improves the robustness in feature generation and evaluation in FELA. For the ablation without the UCB-based learning algorithm, it can be found that the target AUC monotonically increases at a slower pace and converges to a lower value. Hence, proposed UCB-based learning can better balance the exploitation and exploration, which enhances the efficiency of the proposed FELA system.
Figure 11: Ablation study in FELA. The cumulative max AUC during the iterations is tested on the Tencent dataset.
VI Conclusion
@@
Central to FELA is the proposed agentic evolution algorithm, which combines the strengths of reinforcement learning and genetic algorithms to balance exploration and exploitation in a large and heterogeneous feature space. This algorithm, coupled with a hierarchical knowledge base and long–short-term memory design, allows FELA to continuously improve its reasoning, adapt to diverse data domains, and maintain strong explainability throughout the evolution process.
Our experimental and deployment results collectively highlight the potential of LLM-based multi-agent systems in advancing automatic feature engineering. Future work will extend FELA toward broader applications, including multimodal data, dynamic environments, and tighter human-in-the-loop collaboration to further enhance controllability, scalability, and domain alignment.

### v3

@@ V-A Experiment Overview
  • RQ5 (Sensitivity): How sensitive is FELA over the LLM models and hyper-parameter setting?
@@ V-B1 Datasets and Feature Engineering Task
In the experiment, three real-world datasets are adopted for evaluation, including Diabetes Health Indicator Dataset (Dia) [52], Tabao Conversion Prediction Data (Taobao) [2], and the User Churn Data in Tencent Game Platform (Tencent). The basic information of the three datasets is summarized in Table II, where Dia is standard tabular data while the others are event log data with real industrial complexity.
@@
  • Taobao Dataset: This industrial log dataset contains 176,773 users, 9599 items, and 420,721 ad records over 7 days, where each record is labeled with a binary ad impression. The data is organized across five relational tables containing basic identifiers, item profiles, user demographics, contextual signals, and shop reputation metrics, which constitute 31 features for each record. The feature engineering functions jointly process the entire log data and output the features for each user-ad pair. Following standard temporal evaluation protocols, the dataset is chronologically split with the initial 6 days for training and the final day for testing, effectively simulating real-world conversion prediction scenarios where models forecast future conversion behavior based on learned historical patterns.  [^3] The data preprocessing code and initial ideas are released at https://github.com/DKhaoyu/FELA-Feature-Engineering.
  • Tencent Dataset: This dataset is collected from a Massively Multiplayer Online (MMO) game in Tencent for a few days, where the players are sampled from the same user profile to satisfy the i.i.d assumption. Each player is associated with a churn label, indicating whether the user will churn in the next day. In our scenario, each player can have multiple game plays in a day, which results in multiple event log instances. Furthermore, each player in the same game can have multiple teammates, and if the player has more than one teammate, we will separate the records into different rows. Therefore, the instances in such event log data do not follow the i.i.d assumption as mentioned in the introduction session. The feature engineering functions jointly process the entire log data and output the features for each player. The downstream churn prediction is thus working on the player level. For evaluation, a temporal split is adopted: the former day of data are used for training, and the logs from the latter day are held out for testing.
@@ V-B2 Baselines
We evaluate the performance of our proposed FELA system against the following baselines: (1) raw features, (2) OpenFE [60], (3) skope-rules [48], (4) CAAFE [15], (5) FeatLLM [14], (6) AIDE [21], (7) LLM-FE [1]. Here, skope-rules is a traditional rule induction method, the OpenFE is the latest baseline in conventional AutoFE, while CAFFE, FeatLLM, AIDE, and LLM-FE are state-of-the-art LLM-based free coding methods. To ensure a fair comparison, all LLM-based methods, including the baselines and our proposed FELA, utilize the DeepSeek-V3 model via the DeepSeek API by default [29].
@@ V-B3 Metrics
The feature engineering for FELA and the LLM-FE baseline is optimized using the AUC-score, which we denote as the target metric $\mathcal{R}$ in Equation (1). The generated features of all methods are then evaluated using a Random Forest classifier by default. The performance is reported across the main AUC metric and four auxiliary metrics, including Accuracy, Precision, Recall, and F1-score.
TABLE II: Performance comparison on different datasets.
Dataset Method Acc Prec Recall F1 AUC
Dia Raw Features 0.859 0.489 0.166 0.247 0.797
Skope-rules 0.858 0.482 0.171 0.253 0.797
OpenFE 0.858 0.484 0.185 0.268 0.795
CAAFE 0.860 0.499 0.170 0.254 0.799
AIDE 0.859 0.489 0.166 0.247 0.797
@@
Taobao Raw Features 0.630 0.0269 0.580 0.0514 0.630
Skope-rules 0.660 0.0279 0.551 0.0531 0.639
OpenFE 0.592 0.0264 0.630 0.0506 0.636
CAAFE 0.554 0.0243 0.638 0.0471 0.622
AIDE 0.630 0.0264 0.569 0.0504 0.628
FeatLLM 0.645 0.0260 0.536 0.0496 0.631
@@
Tencent Raw Features 0.668 0.576 0.385 0.461 0.683
Skope-rules 0.674 0.595 0.367 0.454 0.685
OpenFE 0.672 0.578 0.418 0.485 0.686
CAAFE 0.671 0.581 0.394 0.470 0.686
AIDE 0.668 0.576 0.385 0.461 0.683
@@
V-C Feature Engineering Performance (RQ1)
Table II presents the feature engineering performance of FELA and baseline methods on different datasets. In the Dia dataset, the proposed FELA achieves superior performance in accuracy, precision, and AUC, demonstrating its effectiveness in handling structured data. Notably, FELA improves the AUC from 0.802 to 0.812 compared to the runner-up method LLM-FE, indicating its capability to generate high-quality discriminative features for tabular data analysis. On the Taobao conversion prediction dataset, FELA maintains its competitive advantage. Our method obtains a notable AUC improvement from 0.641 to 0.653 over LLM-FE. This consistent superiority across multiple metrics confirms FELA’s robustness in processing complex industrial log data with relational characteristics. The evaluation on the Tencent user churn dataset further validates FELA’s superiority on complex industrial log data. FELA outperforms all baselines, achieving a significant AUC boost from 0.686 to 0.701 compared to LLM-FE and OpenFE. These results across three distinct domains demonstrate that FELA consistently generates semantically meaningful features that enhance predictive performance in diverse real-world industrial scenarios.
@@ V-D Steerable Feature Generation (RQ2)
We begin by analyzing the complexity of the generated feature engineering code on the Taobao dataset, as illustrated in Figure 7. The proposed FELA framework generates 22 features derived from 5 distinct feature ideas, achieving an AUC score of 0.653, an improvement of 0.023 over the raw features. This demonstrates FELA’s capability to efficiently produce a substantial number of high-quality features. In contrast, the FeatLLM baseline generates 20 new features, yet yields negligible AUC improvement compared to raw features. Meanwhile, LLM-FE produces 110 code snippets, with the best-performing code improving AUC by 0.011. However, the optimal LLM-FE code only creates only 1 new feature and lacks a continuous evolution mechanism, ultimately limiting its performance potential. Note that the primary requirement for LLM-agent-based feature engineering is the self-evolution capability with elapsed time [5]. The proposed FELA can achieve effective long-horizon self-evolution, which is the key to achieving significant AUC performance improvement. In contrast, the baselines fail to continuously evolve even when increasing the elapsed time.
@@
We further investigate the feature correlation patterns in FELA by analyzing the Pearson correlation heatmap of the aforementioned 22 features, as visualized in Figure 11. The analysis reveals that features derived from the same idea exhibit stronger intra-idea correlation. This phenomenon aligns with expectations, as features originating from the same semantic insight naturally share underlying conceptual foundations. Conversely, inter-idea feature correlations remain relatively low, indicating that FELA successfully explores diverse regions of the feature space. This structured exploration strategy demonstrates FELA’s capability to conduct steerable feature generation while maintaining semantic interpretability. The framework effectively balances between exploiting productive semantic directions and exploring novel feature combinations, resulting in both discriminative power and explanatory value. In Figure 11, we visualize the sampled knowledge database after evolution on the Taobao dataset. It is a super set of the selected features. Green denotes synthesized ideas. Purple denotes new ideas. Red denotes original ideas. Light blue denotes derived features. Thus, we can easily trace the evolution of the knowledge base structure, as well as understand the evolutionary trajectories between the ideas and associated features.
Figure 8: Correlations of selected ideas and features in FELA.
Figure 9: A snapshot of the knowledge database during evolution.
Figure 10: Ablation study of FELA evaluated on the Tencent dataset.
Figure 11: Sensitivity study of UCB-based exploration constant c in FELA
@@ V-E Compatibility to Various Machine Learning Models (RQ3)
TABLE III: AUC performance comparison with different ML models
Method/Model Random Forest XGBoost MLP LR
Raw features 0.683 0.648 0.666 0.650
Skope-rules 0.685 0.647 0.556 0.653
OpenFE 0.686 0.655 0.501 0.650
CAAFE 0.686 0.650 0.558 0.650
AIDE 0.683 0.648 0.666 0.650
FeatLLM 0.657 0.632 0.658 0.646
LLM-FE 0.686 0.653 0.674 0.656
FELA 0.701 0.675 0.673 0.670
We investigate the applicability of the proposed FELA over various machine learning models, including Random Forest, XGBoost, MLP, and Logistic classifiers. Then, the AUC performance of the proposed FELA and the baselines over the Tencent dataset is shown in Table III. It can be observed that the proposed FELA can achieve the best AUC performance over most of the classifiers, where a 0.012 AUC improvement can be achieved on average. Thus, the proposed FELA system can serve as a robust feature engineering solution for various machine learning models.
@@ V-F Abalation Study (RQ4)
We conduct the ablation study in FELA systems to investigate the contribution of its components, including the critic agent, long-short-term memory, and the UCB-based learning algorithm. Then, the AUC in different iterations for the proposed FELA and its ablations is plotted in Figure 11. The ablations without the critic agent and the long-short term memory lead to an obvious AUC decline of around 0.01. Thus, the proposed FELA can benefit from long-short-term memory to exploit the evolutionary experience. Additionally, the critic agent can enhance the correctness and consistency of the generated idea and code, which improves the robustness in feature generation and evaluation in FELA. For the ablation without the UCB-based learning algorithm, it can be found that the target AUC monotonically increases at a slower pace and converges to a lower value. Hence, the proposed UCB-based learning can better balance the exploitation and exploration, which enhances the efficiency of the proposed FELA system.
V-G Sensitivity Study (RQ5)
TABLE IV: Performance of FELA over the Taobao dataset under different LLM models
LLM model Acc Prec Recall F1 AUC
Deepseek-V3 0.678 0.0294 0.551 0.0558 0.653
Kimi-K2-Instruct 0.684 0.0287 0.526 0.0544 0.651
GPT-4o 0.689 0.0294 0.530 0.0556 0.651
We begin by evaluating the sensitivity of FELA to different LLM APIs. Employing the APIs of Kimi-K2 [51] and GPT-4o [20], we report the performance on the Taobao dataset in Table IV. The results show that FELA consistently boosts the AUC to over 0.65 across all backends, improving it by more than 0.02 over the raw features. This consistent gain demonstrates that the performance improvement is attributable to FELA’s novel architecture, rather than being an artifact of a specific LLM’s capabilities.
Next, we further consider the sensitivity of FELA over the exploration constant $c$ in the Tencent dataset, which is shown in Fig. 11. It reveals a unimodal relationship between the exploration constant $c$ and AUC performance, reflecting the classic exploration-exploitation trade-off. Below the optimal value ($c<0.005$), FELA over-exploits operators with high historical scores $\sum_{j}s(d_{i,j})/Q_{i}$, leading to premature convergence. Conversely, when $c>0.005$, excessive exploration of under-sampled operators impedes the refinement of promising candidates. The peak at $c\approx 0.005$ thus represents the ideal balance, maximizing search efficiency and final model performance. Note that proposed FELA can consistently surpass the second optimal baseline Feat-LLM and Open-FE under different $c$, which further validates the system robustness.
VI Conclusion & Future Work
@@
Beyond its academic contributions, FELA has also demonstrated substantial real-world impact. The system has been deployed across several internal data-centric applications to enhance business intelligence and operational efficiency. A representative case involves fraud detection in an MMO game with millions of players, where the goal was to improve the accuracy of a scorecard model through richer player-level features. The baseline includes 60+ expert-crafted features developed by senior data scientists over weeks based on deep domain knowledge. In contrast, we let FELA self-evolve for around 24 hours and keep top 20+ additional features. By incorporating features derived from FELA, the fraud detection system achieved a 47.1% reduction in false positive rate (FPR) compared to expert-engineered baselines—representing a remarkable improvement with significant commercial value.
In future research, we will conduct a more rigorous quantitative evaluation of interpretability. First, we will develop a suite of metrics—such as a concept alignment score to measure the stability of feature importance rankings and counterfactual explanation fidelity to assess local linearity—to objectively compare the explainability of FELA-generated features against baseline methods. Second, we will run a controlled human user study with domain experts to quantitatively assess the comprehensibility, usefulness, and actionability of generated features, providing multi-faceted evidence for the superior human-interpretability of our approach. Finally, we plan to explore a hybrid optimization strategy that combines LLM-driven logical structure generation with efficient numerical methods (e.g., coordinate descent) for parameter tuning, further bridging the gap between automated feature engineering and full-scale AutoML and improving both performance and scalability.